Hello,

I need help finding two consecutive even integers whose product is 224.

Thank you

Answer by scott8148(5891)

consecutive even integers are 2 numbers apart so x(x+2)=224 ... x^2+2x-224=0

factoring gives (x+16)(x-14)=0 ... x+16=0 ... x=-16 ... x-14=0 ... x=14

the two numbers are 14 and 16; but they can both be positive or both be negative